# High-Quality Image Restoration Following Human Instructions

### 29 Jan 2024 v1

@@ Abstract
Figure 1: Given an image and a prompt for how to improve that image, our all-in-one restoration model corrects the image considering the human instruction. InstructIR, can tackle various types and levels of degradation, and it is able to generalize to restoring real photos.
@@ 2 Related Work
Figure 2: We train our blind image restoration models using common image datasets, and prompts generated using GPT-4, note that this is (self-)supervised learning. At inference time, our model generalizes to human-written instructions and restores real images.
@@ 3.2 Text Encoder
Figure 3 shows that while the text encoder is capable out-of-the-box to cluster instructions to some extent (Figure 2(a)), our trained projection yields greatly improved clusters (Figure 2(b)). We distinguish clearly the clusters for deraining, denoising, dehazing, deblurring, and low-light image enhancement. The instructions for such tasks or degradations are very characteristic. Furthermore, we can appreciate that “super-res” and “enhancement” tasks are quite spread and between the previous ones, which matches the language logic. For instance “add details to this image” could be used for enhancement, deblurring or denosising. In our experiments, $d_{t}\!=\!384$, $d_{v}\!=\!256$ and $\mathbf{W}$ is a linear layer.
@@
We propose a guidance loss on the text embedding $\mathbf{e}$ to improve training and interpretability. Using the degradation types as targets, we train a simple classification head $\mathcal{C}$ such that $\mathbf{c}=\mathcal{C}(\mathbf{e})$, where $\mathbf{c}\in\mathrm{R}^{D}$, being $D$ is the number of degradations. In our experiments $D=7$ and the classification head $\mathcal{C}$ is a simple two-layers MLP. Thus, we only need to train a projection layer and a simple MLP to capture the natural language knowledge. This allows the text model to learn meaningful embeddings as we can appreciate in Figure 3, not just guidance vectors for the main image processing model.
We find that the model is able to classify accurately (i.e. over 95% accuracy) the underlying degradation in the user’s prompt after a few epochs.
@@ 4 Experimental Results
We provide extensive qualitative results using benchmark images in Figures 1718 and 19. We also evaluate our model on 9 well-known benchmarks for different image restoration tasks: image denoising, deblurring, deraining, dehazing, and image enhancement. We present extensive quantitative results in Table 2. Our single model successfully restores images considering different degradation types and levels. We provide additional results and ablation studies in the supplementary material.
@@ 4.2 Datasets and Benchmarks
Table 2: Quantitative results on five restoration tasks (5D) with state-of-the-arts general image restoration and all-in-one methods. We highlight the reference model without text, the best results, and the second best results. We also present the ablation study of our multi-task variants (from 5 to 7 tasks — 5D, 6D, 7D). This table is based on Zhang et al. IDR [100] (CVPR ’23).
@@ 7 Conclusion
We define our loss functions in the paper Sec. 4.1. Our training loss function is $\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{ce}$, which includes the loss function of the image model ($\mathcal{L}_{1}$), and the loss function for intent (task/degradation) classification ($\mathcal{L}_{ce}$) given the prompt embedding. We provide the loss evolution plots in Figures 14 and 15. In particular, in Figure 15 we can observe how the intent classification loss (i.e. predicting the task (or degradation) given the prompt), tends to 0 very fast, indicating that our language model component can infer easily the task given the instruction.
@@
We present diverse qualitative samples in Figures 1718, and 19. Our method produces high-quality results given images with any of the studied degradations. In most cases the results are better than the reference all-in-one model AirNet [42]. Download all the test results at https://github.com/mv-lab/InstructIR.

### 21 Feb 2024 v3

# InstructIR: High-Quality Image Restoration Following Human Instructions

@@ Abstract
Figure 1: Given an image and a prompt for how to improve that image, our all-in-one restoration model corrects the image considering the human instruction. InstructIR, can tackle various types and levels of degradation, and it is able to generalize in some real-world scenarios.
@@ 2 Related Work
Figure 2: We train our blind image restoration models using common image datasets, and prompts generated using GPT-4, note that this is (self-)supervised learning. At inference time, our model generalizes to human-written instructions and restores (or enhances) the images.
@@ 3.2 Text Encoder
Figure 3 shows that while the text encoder is capable out-of-the-box to cluster instructions to some extent (Figure 2(a)), our trained projection yields greatly improved clusters (Figure 2(b)). We distinguish clearly the clusters for deraining, denoising, dehazing, deblurring, and low-light image enhancement. The instructions for such tasks or degradations are very characteristic. Furthermore, we can appreciate that “super-res” and “enhancement” tasks are quite spread and between the previous ones, which matches the language logic. For instance “add details to this image” could be used for enhancement, deblurring or denosising. In our experiments, $d_{t}\!=\!384$, $d_{v}\!=\!256$ and $\mathbf{W}$ is a linear layer. The representation $\mathbf{e}$ from the text encoder is shared across the blocks, and each block has a trainable projection $\mathbf{W}$.
@@
We propose a guidance loss on the text embedding $\mathbf{e}$ to improve training and interpretability. Using the degradation types as targets, we train a simple classification head $\mathcal{C}$ such that $\mathbf{c}=\mathcal{C}(\mathbf{e})$, where $\mathbf{c}\in\mathrm{R}^{D}$, being $D$ is the number of degradation classes. The classification head $\mathcal{C}$ is a simple two-layers MLP. Thus, we only need to train a projection layer $\mathbf{W}$ and a simple MLP to capture the natural language knowledge. This allows the text model to learn meaningful embeddings as we can appreciate in Figure 3, not just guidance vectors for the main image processing model. We find that the model is able to classify accurately (i.e. over 95% accuracy) the underlying degradation in the user’s prompt after a few epochs.
@@ 4 Experimental Results
We provide extensive qualitative results using benchmark images in Figures 1920 and 21. We also evaluate our model on 9 well-known benchmarks for different image restoration tasks: image denoising, deblurring, deraining, dehazing, and image enhancement. We present extensive quantitative results in Table 2. Our single model successfully restores images considering different degradation types and levels. We provide additional results and ablation studies in the supplementary material.
@@ 4.2 Datasets and Benchmarks
Table 2: Quantitative results on five restoration tasks (5D) with state-of-the-art general image restoration and all-in-one methods. We highlight the reference model without text (image only), the best overall results, and the second best results. We also present the ablation study of our multi-task variants (from 5 to 7 tasks — 5D, 6D, 7D). This table is based on Zhang et al. IDR [100] (CVPR ’23).
@@ 7 Conclusion
Instruction: “my colors are too off, make it pop so I can use these photos in instagram”
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Input InstructIR (Output)
Figure 14: Image Enhancement Results. We provide qualitative samples from the MIT5K Dataset [5].
Instruction: “the image is too dark, it has poor illumination, can you make it brighter?”
Refer to caption Refer to caption
Refer to caption Refer to caption
Instruction: “Please, reduce the motion in this image so it is more clear”
Refer to caption Refer to caption
Figure 15: Additional high-resolution qualitative results using the LOL [83] dataset (low-light image enhancement), and the GoPro [57] dataset (motion deblurring). We provide the corresponding natural language instructions.
@@
We define our loss functions in the paper Sec. 4.1. Our training loss function is $\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{ce}$, which includes the loss function of the image model ($\mathcal{L}_{1}$), and the loss function for intent (task/degradation) classification ($\mathcal{L}_{ce}$) given the prompt embedding. We provide the loss evolution plots in Figures 16 and 17. In particular, in Figure 17 we can observe how the intent classification loss (i.e. predicting the task (or degradation) given the prompt), tends to 0 very fast, indicating that our language model component can infer easily the task given the instruction.
@@
We present diverse qualitative samples in Figures 1920, and 21. Our method produces high-quality results given images with any of the studied degradations. In most cases the results are better than the reference all-in-one model AirNet [42]. Download all the test results at https://github.com/mv-lab/InstructIR.